# SIM2VR: Towards Automated Biomechanical Testing in VR

### v1

@@ 5. Predicting Interactive User Behaviour with SIM2VR
In this section, we show how SIM2VR can be used to predict movement-based interaction in VR and demonstrate its utility in evaluating VR interface design choices. We developed a VR version of the popular, visuomotorically demanding, fast-paced and dynamic arcade game Whac-A-Mole in Unity with different interface variations. With SIM2VR, we train a UitB model to play the game, and then use this model to predict differences in performance, effort and strategies between different game variations. To assess the quality of these predictions, we compare them to reference data from a user study, where real humans were asked to play the same VR game.
We re-implemented our own version of the Whac-A-Mole game, as this allows us to investigate in a controlled way the extent to which SIM2VR can predict the effect of design choices such as difficulty or target area placement. In the supplementary material, we show how to integrate SIM2VR into an existing (third-party) Unity application and discuss how to create a simulated user in UitB.
5.1. Game Design
@@
In the Whac-A-Mole game the goal is to hit targets (the ”moles”) with a hammer to score points. The targets appear randomly on a $3\times 3$ grid as shown in Figure 6a. If a target is successfully hit within one second, it explodes and disappears (see Figure 2) and the game score is increased by one. Otherwise, the target collapses after one second and no points are given. One round of game play lasts for one minute.
@@ 5.2. Training Simulated Users
Figure 7. Still frames from an evaluation of a simulated user as it is hitting a target on the right side of the target area. The top row shows the simulated user in MuJoCo, while the bottom row is a video capture from the Unity application.
@@
We train a UitB user model to play the different variations of the game. We use MoblArmsWrist, a MuJoCo version of the MoBL ARMS model (Saul et al., 2014) of the right arm, as the biomechanical model. This model includes shoulder, elbow and wrist joints, and the muscles actuating these joints, resulting in a total of 7 DOFs and 32 muscles. Since the model does not actively model neck and head movements, the position and orientation of the HMD were kept constant throughout the simulation. We use the UnityHeadset vision module provided by SIM2VR, along with the pre-implemented UitB proprioception module, to generate observations during training. The UnityHeadset module is configured to include the green colour and depth channels, without stacking visual observations. The proprioception module provides observations regarding the biomechanical model, such as the joint positions, velocities, and accelerations, as well as internal muscle activation states.
For training, we use a reward function consisting of multiple components:
(1) $r(s,a)=w_{s}*S+w_{c}*v_{h}*C_{c}+w_{d}*C_{d}+w_{e}*C_{e}$
The component $S$ is the game score provided by the game; $C_{c}$ is a term to penalize target contacts with too low a hitting velocity in the constrained version (to encourage faster hits); $C_{d}$ penalises the distance between the hammer and each active target (to incentivise movement toward any target); and $C_{e}$ is an effort term based on the 3CC-r fatigue model (Looft et al., 2018; Cheema et al., 2020). The reward components are weighted with $w_{s}=10$, $w_{c}=2.5$, $w_{d}=1$ and $w_{e}=0.1$. The target contact term $C_{c}$ is additionally weighted by the current hammer velocity $v_{h}$.
Simulated users were trained separately for the constrained and unconstrained versions, while keeping the difficulty level fixed to medium and sampling the target area placements randomly at the beginning of each episode. [^10] While it may be reasonable to continue training the user models on either the easy or hard difficulty levels, we observed that the models trained solely on the medium level already generalise well to the remaining difficulties. For each version we train two instances of the simulated user, called SIMu and SIMa. SIMu is trained on the default game dynamics, i.e., the position of the ”mole” is sampled uniformly randomly among the nine possible positions. SIMa, on the other hand, is trained using an adaptive automated curriculum manually implemented in Unity. Such a curriculum can be thought of as a “personal trainer” that creates a customised training plan for the simulated user based on its current strengths and weaknesses.
@@ 5.3. User Study Method
In order to compare the simulated user behaviour against real player behaviour, we conducted a user study where we collected game playing data from 18 right-handed participants. Twelve participants played the the constrained version of the game, while six participants played the unconstrained version. Each participant played the same six difficulty $\times$ target area placement configurations used to evaluate the simulated users.
@@
The participants were mostly local students (graduate or post-graduate) and faculty members at Aalto University, Finland. The average age was 28.8 years, with a standard deviation of 6.2 years. Eight of the participants identified as female, nine as male, and one as non-binary.
Task
The participants first played the three difficulty levels, while keeping the target area placement fixed to mid, in a counterbalanced order. Participants were then given a rest period of at least 5 minutes, with the option to rest longer if necessary to ensure their right arm was not fatigued. After resting, participants played the game in all three target area placements, in a counterbalanced order, with the difficulty level fixed to medium, and reported the Borg Rating of Perceived Exertion (Borg RPE) (Borg, 1982) score after each round.
@@
The experiments were conducted in an office room. The participants were seated, but not restricted to the chair. Participants were asked to try to avoid bending and rotating their torso (as the simulated users have a fixed torso), and focus on their arm movements. Participants were allowed to train with the medium difficulty level with mid target area placement. This was to reduce learning effects, and to provide the participants a chance to familiarise themselves with the notion of hitting non-physical floating targets. The experiment length was approximately 30 minutes, and participants were rewarded with a 15€ gift card to a local restaurant.
@@ 5.4.1. Performance
Figure 8. The simulated users SIMu and SIMa (left and mid bars) show similar numbers of target hits and misses in the Whac-A-Mole game as the 12 real users from our study on average (right bars). Differences between the three difficulty levels are predicted well.
@@
Figure 9. Both SIMu and SIMa predict target hit rates that are within the between-user variability observed from real players. The simulations also show a similar decrease in hit rate as the game difficulty increases as most real players.
@@ 5.4.2. Effort
User Variable Placement Mean/ Median Std./ IQR Alternative Hypothesis Wilcoxon Signed Rank Z-score p-value   SIMu Max. Fatig. MUs Low 0.232342 0.001100 Low ¡ Mid 3.059 1.0 (n.s.) Mid 0.203537 0.001641 Mid ¡ High -3.059 0.0002 (***) High 0.238448 0.001772 Low ¡ High -3.059 0.0002 (***) SIMa Max. Fatig. MUs Low 0.265645 0.002151 Low ¡ Mid -2.51 0.0046 (**) Mid 0.269120 0.002360 Mid ¡ High 3.059 1.0 (n.s.) High 0.229852 0.001937 Low ¡ High 3.059 1.0 (n.s.) User Study Borg RPE Low 9 1.25 Low ¡ Mid -1.508 0.0658 (n.s.) Mid 9 3 Mid ¡ High -2.414 0.0079 (**) High 9.5 4.25 Low ¡ High -2.213 0.0134 (*)
Table 2. Overview of both descriptive and inferential statistics for effort predictions. Descriptive statistics include the mean and standard deviation for the maximum fatigued motor units per round (averaged over all muscles), as predicted by the simulation, and the median and interquartile range (IQR) for the Borg RPE as reported in the user study ()). Kolmogorov-Smirnov tests showed that none of the considered variables are normally distributed (all p-values ¡ 0.001); we thus used one-sided Wilcoxon Signed rank tests to infer significant differences between the three target area placements.
Explanations: *: $p\leq 0.05$; **: $p\leq 0.01$; ***: $p\leq 0.001$; n.s.: $p>0.05$.
Abbreviations: Std. for standard deviation, IQR for interquartile range, Max. Fatig. MUs for maximum fatigued motor units and n.s. for not significant.
[…]
The table consists of three parts. The left part has the columns ”User” and ”Variable” and the following rows: 1. SIM_u - Max. Fatig. MUs. 2. SIM_a - Max. Fatig. MUs. 3. User Study - Borg RPE. The second part, which is placed right next to the first part, has the columns ”Placement”, ”Mean/Median †” and ”Std./IQR †” and the following rows (three subrows per row of the first part): 1a. Low - 0.232342 - 0.001100. 1b. Mid - 0.203537 - 0.001641. 1c. High - 0.238448 - 0.001772. 2a. Low - 0.265645 - 0.002151. 2b. Mid - 0.269120 - 0.002360. 2c. High - 0.229852 - 0.001937. 3a. Low - 9 † - 1.25 †. 3b. Mid - 9 † - 3 †. 3c. High - 9.5 † - 4.25 †. The third part, which is placed right next to the first second, has the columns ”Alternative Hypothesis”, ”Wilcoxon Signed Rank – Z-score” and ”Wilcoxon Signed Rank – p-value” and the following rows (three subrows per row of the first part): 1a. Low ¡ Mid - 3.059 - 1.0 (n.s.). 1b. Mid ¡ High - -3.059 - 0.0002 (***). 1c. Low ¡ High - -3.059 - 0.0002 (***). 2a. Low ¡ Mid - -2.51 - 0.0046 (**). 2b. Mid ¡ High - 3.059 - 1.0 (n.s.). 2c. Low ¡ High - 3.059 - 1.0 (n.s.). 3a. Low ¡ Mid - -1.508 - 0.0658 (n.s.). 3b. Mid ¡ High - -2.414 - 0.0079 (**). 3c. Low ¡ High - -2.213 - 0.0134 (*). In the ”Wilcoxon Signed Rank – p-value” column, all cells with at least one ”*” (i.e., those showing significant differences) are highlighted in green.
To assess the impact of target area placement on movement effort, participants in the user study were asked to report their perceived exertion in terms of the Borg RPE scale, after playing each 1-minute round. The Borg RPE reported for the high placement was significantly higher than for both the low and mid placements, while no significant difference could be found between low and mid placements (summary statistics as well as details on the statistical testing are given in Table 2). This also agrees well with qualitative remarks of most of the participants, stating that the arm felt considerably more fatigued after playing the high placement as compared to the low and mid placements. To estimate the level of exertion from our simulation, we measured the percentage of fatigued motor units as predicted by the 3CC-r fatigue model for each simulation time frame, calculated the maximum value per round for each muscle separately, and then took the average over all 32 muscles.
As shown in Table 2, SIMu predicts a significant increase in fatigue between the mid and high as well as between the low and high placements, showing the same trend as observed for the Borg RPE scale for real players. Interestingly, the fatigue predicted for the low placement is considerably higher than in mid placement. Comparing the simulation videos for these placements, we found that in the low placement SIMu follows a strategy that hardly flexes the elbow, but rather keeps the arm extended for almost the entire movement, which may lead to higher muscle fatigue over time.
For SIMa, the mid placement exhibits significantly higher fatigue values than the low placement ($p=0.0046$), whereas no significant increase from mid to high or from low to high could be found. We found that this is mainly due to different strategies for each of the three placements, despite being trained together and sharing the same network weights. In particular, SIMa has learned to efficiently reduce unnecessary shoulder movements mainly in the high placement, resulting in a lower number of fatigued motor units than in the low and mid placements.
Our findings can be seen as an example of how small changes in the training may have a considerable impact on which strategies are learned and predicted by the user simulation. While we demonstrate that characteristic differences in subjective levels of muscle exertion can be predicted in principle by our simulation-based approach, effort-related predictions must currently be treated with caution, as the effort can vary with small changes in the movement strategy.
5.4.3. Strategy
Figure 10. In the unconstrained variant our simulation predicts a different strategy, with the hammer kept close to the target plane during the entire movement (orange line in left plot). This strategy was also observed among real players (orange line in right plot). The hammer depth trajectories of (different) simulated and real players interacting with the constrained version are shown for comparison (blue lines). All movements are in the hard level, but similar results were also observed for the medium level.
[…]
Left (Subfigure title: ”Simulated Users – Hammer Depth During Movement”): Hammer depth in meters (y-axis) plotted over time (x-axis) for an entire round of 60 seconds, once for a simulated user movement in the constrained variant (blue line) and once in the unconstrained variant (orange line). Right (Subfigure title: ”Real Users – Hammer Depth During Movement”): Hammer depth in meters (y-axis) plotted over time (x-axis) for an entire round of 60 seconds, once for a real user movement in the constrained variant (blue line) and once in the unconstrained variant (orange line).
Simulations trained via SIM2VR can anticipate the effect of a simple game design choice on the strategies employed by users. In Figure 10 this is shown for a trajectory comparison between the constrained and the unconstrained version of the game. In the unconstrained version, an entirely different strategy is predicted by our simulation; in particular, the unconstrained policy retracts the arm considerably less, and instead keeps the hammer relatively close to the target area during movements, which indeed allows to hit many more targets in the same time. [^11] Applying the two simulation strategies to the hard level, 195 target hits were observed in the unconstrained variant, and 102 target hits and 35 additional target contacts with too low hitting velocity were observed in the constrained variant. Interestingly, one of the six users who played the unconstrained version of the game exploited the same strategy, as shown in Figure 10 (right, orange line) along with a reference user trajectory from the constrained version of the game (blue line). [^12] Another user pursued a similar strategy but hit the targets more from above, while the remaining four users followed essentially the same strategy as the users in the constrained version.
This demonstrates the value of biomechanical user models in discovering specific strategies that users may exploit for a given game design. By making such insights available early in the VR development process, SIM2VR can help find game dynamics that meet the designers’ requirements for performance, effort, comfort, space limitations and more.
5.5. Summary
This demonstration shows that SIM2VR can be utilised to predict differences in terms of performance, effort and strategies between different versions of a VR game. Our approach enables the evaluation and comparison of different prototypes and design choices, and thus could be used in addition to (or as a partial replacement for) user studies in the early stages of application development.
6. Discussion
Computational approaches to modelling and simulation of users have the potential to be of value in design and analysis of interactive systems. However, in addition to the challenges of modelling the complexity of human behaviour, a significant challenge is modelling the sensory experience of users, who can be in complex environments and subject to unmodelled external disturbances to all senses, which can reduce the potential benefits of a “sim-to-real” transfer (a term mainly used in the field of robotics (Dimitropoulos et al., 2022)).
VR interaction, however, is a special case of human–computer interaction in that significant aspects of the user’s perception (namely vision and audio) can be controlled by the VR application to a greater extent than other cases. In particular, the (currently) low number of different VR systems, the “full-screen mode” of most VR applications and the typically high level of immersion (Seinfeld et al., 2020) reduces the impact of external disturbances when assessing performance. VR interaction thus offers ideal conditions for creating ecologically valid simulations and performing automated user testing. This makes VR a good starting point for research into simulation models which have well-aligned information from visual and auditory displays.
With SIM2VR, we have built on these ideal conditions and developed a system that does not only give a biomechanical user model access to the same VR application as real users, but also ensures that it can interact with the application just as if a physical VR controller and HMD were used. This constitutes a decisive step towards automated biomechanical testing in VR. In the future, biomechanical user simulations could lead to more comfortable and health-promoting VR interactions, as well as VR environments that specifically consider the abilities of individual user groups.
6.1. Scope and Limitations
SIM2VR mainly addresses the application-related dimensions of the alignment problem. Aligning the user-related dimensions, i.e., physics, perception, cognition, however, remains challenging. This particularly includes finding an appropriate biomechanical user model, modelling multisensory noisy perception, and designing a task-specific reward function, as we discuss in the following.
Due to its dependency on UitB, SIM2VR currently only provides the standard MoBL ARMS model (Saul et al., 2014), which may be unsuitable for simulating dexterous hand movements or body rotations. However, more sophisticated models of the human body focusing on either the elbow, hand, single finger or lower extremity have recently been released as part of the MyoSuite project [^13] https://sites.google.com/view/myosuite and can be integrated directly into UitB. In addition, the MyoConverter (Ikkala and Hämäläinen, 2020; Wang et al., 2022) can be used to transfer OpenSim models to MuJoCo.
Regarding the interaction technique, in this paper we have focused on one-handed VR interaction using a standard VR controller and HMD. An extension of SIM2VR to other movement-based methods such as full body movements sensed via Motion Capture is straightforward. The same applies to currently missing output modalities such as auditory or haptic feedback.
The vision module provided by SIM2VR currently only models the user’s visual perception of the virtual environment at a very rudimentary level. To better reflect the complexity and subtlety of visuomotor VR interaction, this module should be extended in the future to include, for example, binocular, foveal, noisy and/or delayed visual observations. Initial ideas for modelling these well-established phenomena might be derived from (Strasburger et al., 2011; Kowler, 2011; Fischer et al., 2022).
Simulating user cognition is known to be particularly challenging. In the context of reinforcement learning, a major part of cognition modelling is to identify appropriate reward terms that incentivise “human-like” behaviour. However, no general guidelines have yet been proposed on how to define these reward terms appropriately for a given interaction task. Additionally, reward functions can easily induce unintended behaviour, if not designed carefully. For example, a reward signal that penalises the remaining distance to multiple positions at the same time (e.g., to all currently active targets) will encourage an ”intermediate” strategy that opts for a particular target position later rather than sooner. Such a “strategic bias” may also degrade the quality of simulation-based predictions. The availability of highly skilled and approved simulated users for a variety of interaction tasks will therefore be critical to the extent to which biomechanical simulation finds its way into future automated VR testing routines.
In order to expand the scope of automated biomechanical testing beyond simple visuomotor interaction tasks, further research into (and validation of) more complex RL-based training methods and design choices of reward functions and metaparameters is certainly needed. Promising RL methods and frameworks in this regard include Hierarchical RL (Li et al., 2019; Nachum et al., 2018), in which a complex learning problem is decomposed into simpler subtasks that are easier to learn, and Transfer Learning (Zhu et al., 2020), in which models are first exposed to a data-rich domain before being fine-tuned in a data-scarce or otherwise more challenging environment. Moreover, recent successes in training user models for cognitively and biomechanically challenging motor control and interaction tasks, such as rotating multiple Baoding balls over the palm of the hand (Caggiano et al., 2022b) or parking a remote-controlled car using a joystick (Ikkala et al., 2022), raise hopes for versatile simulated users that can be integrated directly into a range of existing VR applications without extensive refinement.
7. Conclusion
The SIM2VR system presented in this paper enables the training of RL-based simulated users directly in the same VR environments that real users interact with. By partially addressing the VR Simulation Alignment Problem, our approach increases the ecological validity of biomechanical user simulations in VR. Using Whac-A-Mole as an example, we have demonstrated that SIM2VR can predict characteristic differences between different game variants in terms of performance and effort, and reveal potential user strategies. By lowering the barrier to the use of biomechanical simulation, we hope that SIM2VR will increase the adoption of automated testing methods in VR development. Given that biomechanical user models can help identify bugs, safety risks and health issues prior to user testing, we expect user modeling to assume a position as an indispensable part of the VR development process in the future.

### v2

@@ 5. Predicting Interactive User Behaviour with SIM2VR
In this section, we show how sim2vr can be used to predict movement-based interaction in VR and demonstrate its utility in evaluating VR interface design choices. We therefore developed a VR version of the popular, visuomotorically demanding, fast-paced and dynamic arcade game Whac-A-Mole in Unity with different interface variations. This allows us to investigate in a controlled way the extent to which sim2vr can predict the effect of design choices such as difficulty or target area placement. With sim2vr, we train a UitB model to play the game, and then use this model to predict differences in performance, effort and strategies between different game variations. To assess the quality of these predictions, we compare them to reference data from a user study, where real humans were asked to play the same VR game.
@@
Figure 6. (a) In Whac-A-Mole, targets appear randomly at one of nine fixed positions and must be hit with a hammer within one second to score a point. The transparent targets are shown for visualisation purposes only and are not visible during the game. (b) The game allows for three different placements of the target area.
@@
Top (subcaption: ”Front View of Target Grid”): Front view of the target plane and the target grid in the Whac-A-Mole game, together with the controlled hammer hitting a target. Bottom (subcaption: ”Target Area Placements”): Three target areas (low, mid, high) differing in terms of position and orientation are shown in 3D space together with the controlled hammer.
5.1. Game Design
In the Whac-A-Mole game the goal is to hit targets (the ”moles”) with a hammer to score points. The targets appear randomly on a $3\times 3$ grid as shown in Figure 6a. If a target is successfully hit within one second, it explodes and disappears (see Figure 2) and the game score is increased by one. Otherwise, the target collapses and no points are given. One round of game play lasts for one minute.
@@ 5.2. Training Simulated Users
We train a UitB user model to play the different variations of the game. We use MoblArmsWrist, a MuJoCo version of the MoBL ARMS model (Saul14) of the right arm, as the biomechanical model. This model includes shoulder, elbow and wrist joints, and the muscles actuating these joints, resulting in a total of 7 DOFs and 32 muscles. As the model does not actively model neck and head movements, the pose of the virtual HMD is constant throughout the simulation. To account for this limitation, the target area was designed to be of a sufficiently small size to fit within the field of view of the participants, thus minimising the necessity for head movements. Furthermore, we use a non-scaled version of the biomechanical model, as we are not attempting to replicate a specific participant’s movements, but rather aim to demonstrate sim2vr’s utility in generating movements that could be produced by a human participant. Simulating the between-user variability by simultaneously training differently scaled biomechanical models is left for future work.
We use the UnityHeadset vision module provided by sim2vr, along with the pre-implemented UitB proprioception module, to generate observations during training. The UnityHeadset module is configured to include the green colour and depth channels [^10] As the target colour changes from green to red, it is sufficient to use either the green or red channel to infer whether a target is about to disappear., without stacking visual observations. The proprioception module provides observations regarding the biomechanical model, such as the joint positions, velocities, and accelerations, as well as internal muscle activation states.
For training, we use a reward function consisting of multiple components, which are obtained from states of the VR environment, such as game scores, and from states of the biomechanical model, such as muscle fatigue:
(1) $r(s,a)=w_{s}S+w_{c}v_{h}C_{c}+w_{d}C_{d}+w_{e}C_{e}$
The component $S$ is the game score provided by the game; $C_{c}$ is a term to penalise target contacts with too low a hitting velocity in the constrained version (to encourage faster hits); $C_{d}$ penalises the sum of the distances between the hammer and each active target (to incentivise movement toward any target); and $C_{e}$ is an effort term based on the 3CC-r fatigue model (Looft18; Cheema20). The reward components are weighted with $w_{s}=10$, $w_{c}=2.5$, $w_{d}=1$ and $w_{e}=0.1$. The target contact term $C_{c}$ is additionally weighted by the current hammer velocity $v_{h}$.
Simulated users were trained separately for the constrained and unconstrained versions, while keeping the difficulty level fixed to medium and sampling the target area placements randomly at the beginning of each episode. [^11] While it may be reasonable to continue training the user models on either the easy or hard difficulty levels, we observed that the models trained solely on the medium level already generalise well to the remaining difficulties. For each version we train two instances of the simulated user, called SIMu and SIMa. SIMu is trained on the default game dynamics, i.e., the position of the ”mole” is sampled uniformly randomly among the nine possible positions. SIMa, on the other hand, is trained using an adaptive automated curriculum manually implemented in Unity. Such a curriculum can be thought of as a “personal trainer” that creates a customised training plan for the simulated user based on their current strengths and weaknesses.
Figure 7. Still frames from an evaluation of a simulated user playing Whac-A-Mole. The top row shows the simulated user in MuJoCo, while the bottom row is a video capture from the Unity application. In a typical hitting movement, the simulated user gains momentum (left), strikes the target with the virtual hammer (mid) and moves the hammer away from the target area by bending the elbow (right).
@@ 5.3. User Study Method
In order to compare the simulated user behaviour against real player behaviour, we conducted a user study where we collected game playing data from 18 participants. All but two participants were students and faculty members at Aalto University; one participant was a student at the University of Helsinki, and one participant was not affiliated with either university. All participants were right-handed. The average age was 28.8 years, with a standard deviation of 6.2 years. Eight of the participants identified as female, nine as male, and one as non-binary.
Experimental Design
Twelve participants played the constrained version of the game, while six participants played the unconstrained version. Each participant played the six difficulty $\times$ target area placement configurations that were also used to evaluate the simulated users. For each trial, we recorded the participant’s body and hammer movement. For the three target area placement configurations (with medium difficulty), we also asked participants to report the Borg Rating of Perceived Exertion (Borg RPE) (Borg82) score after each round.
@@
The experiments were conducted in an office room. The participants were seated, but not restricted to the chair. Participants were asked to try to avoid bending and rotating their torso (as the simulated users have a fixed torso), and focus on their arm movements. Participants were allowed to train with the medium difficulty level with mid target area placement. This was to reduce learning effects, and to provide the participants a chance to familiarise themselves with the notion of hitting non-physical floating targets. The participants first played the three difficulty configurations in a counterbalanced order, and were then given a rest period of at least 5 minutes, with the option to rest longer if necessary to ensure their right arm was not fatigued. Finally, participants played the three target area placement configurations in a counterbalanced order, and reported the Borg RPE after each round. The experiment length was approximately 30 minutes, and participants were rewarded with a 15 euro gift card to a local restaurant.
@@ 5.4.1. Performance
Figure 8. The simulated users SIMu and SIMa (left and mid bars) show similar numbers of target hits and misses in the constrained version of the Whac-A-Mole game as the 12 real users from our study on average (right bars). Differences between the three difficulty levels are predicted well.
@@
Figure 9. Both simulated users SIMu and SIMa (highlighted in red) predict target hit rates that are within the between-user variability observed from real players for the constrained version of the Whac-A-Mole game. The simulations also show a similar decrease in hit rate as the game difficulty increases as most real players.